A copper calorimeter that weighs 300g contains 200g Oil initially at 293.15K. 180g of Aluminium at 300°C is added to the oil. What will be the final temperature of the system after equilibrium is established?

Specific heat capacity for Copper = 0.379J/g°C

Specific heat capacity for Aluminium= 0.921J/g°C

Specific heat capacity for Oil= 1.54 J/g°C


1
Expert's answer
2020-11-05T10:10:49-0500

"m_c = 300g\\
m_o = 200g\\
T_c = 293.15K = 20°C\\
T_o = 20°C\\
m_a = 180g\\
T_a = 300°C\\
T_{final} = ?"


"c_c = 0.379\ J/g°C\\
c_a = 0.921\ J/g°C\\
c_o = 1.54\ J/g°C"



According to the Law of conservation of energy,

Heat gained by substance A = Heat lost by substance B


"\rightarrow m_c c_c (T_f - T_c) + m_o c_o (T_f - T_o) = m_a c_a (T_a T_f)\\
\rightarrow 300×0.379(T_f -20) + 200×1.54(T_f-20) = 180×0.921(300-T_f)\\
\rightarrow 113.7(T_f-20) + 308(T_f-20) = 165.78(300-T_f)\\
\rightarrow (113.7+308)(T_f-20)= 49734-165.78T_f\\
\rightarrow 421.7(T_f-20)= 49734-165.78T_f\\
\rightarrow 421.7T_f - 8434=49734-165.78T_f\\
\rightarrow 587.48T_f = 58168\\
\rightarrow T_f = 99.01°C"


"\therefore" The final temperature of the system after equilibrium is established will be "99.01\u00b0C"